dariusthegoat95:

does anyone know radical 8 time -3 radical 2

supie:

\(\sqrt{8} \times -3 \sqrt{2}\) ?

dariusthegoat95:

yea that

snowflake0531:

nvmd, ii wrote it wrong lol

AZ:

or you could just \(\sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\) and in this case, you can easily simplify it

snowflake0531:

lol, ye that ^

snowflake0531:

that would be much easier, multiply the insides and the outsides so it would be -3 times sqrt(8 * 2)

snowflake0531:

which is -3 sqrt16 do you know how to simply that ^

dariusthegoat95:

24 radical 2

dariusthegoat95:

so 12

dariusthegoat95:

i think

snowflake0531:

actually... it would be negative since it would be -3 times 4

AZ:

you can't multiply a number inside the radical with a number outside

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dariusthegoat95 24 radical 2 \(\color{#0cbb34}{\text{End of Quote}}\) btw, how'd you get this...

AZ:

\(\sqrt{8} \times -3 \sqrt{2}\) \( = -3 \sqrt{8 \times 2}\) what is 8 * 2 =

AZ:

He did 3 * 8 @snowflake0531

snowflake0531:

oh- lol oop

AZ:

what is 8 times 2?

dariusthegoat95:

i messed somethig up

dariusthegoat95:

it will be -12 radical 4

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dariusthegoat95 it will be -12 radical 4 \(\color{#0cbb34}{\text{End of Quote}}\) ummm not that either first, what's sqrt 16?

dariusthegoat95:

-3 radical 16 is because 16 is a perfect square of 4 so -3 times 4 is -12 with a radical of 4

snowflake0531:

then there wouldn'tbe a radical anymore

AZ:

no you were on the right tract \(\sqrt{16} = 4\) after that all you have left is -3 * 4

AZ:

there is no more radical after that

dariusthegoat95:

ok thank you

AZ:

No problem!